OpenStudy (yanasidlinskiy):

What is the probability that one is not nervous around strangers? The probability that both will be nervous is 0.0324

OpenStudy (agent47):

One is not nervous = 1 - P(None are nervous) - P(Both are nervous)

OpenStudy (agent47):

I feel like it's missing something.. is there any other info?

OpenStudy (yanasidlinskiy):

About 18​% of the population of a large country is nervous around strangers. If two people are randomly​ selected, what is the probability both are nervous around strangers​? What is the probability at least one is nervous around strangers​?

OpenStudy (yanasidlinskiy):

I got the first part I just need help on the second part. I have no idea where to start and where to end.

OpenStudy (agent47):

oh so there you go, 18% is nervous, so the remaining 82% aren't. P(none are nervous)=0.82 1-.82-.0324

OpenStudy (agent47):

Meaning 1 - P(none are nervous) - P(both are nervous) Have someone double check tho, I'm not too sure about probability

OpenStudy (yanasidlinskiy):

So I would find the remaining and subtract it from 1?

OpenStudy (agent47):

you need to find P that none are nervous, which is 1-.18=.82 P(one out of two is not nervous) = P(only one out of the two is nervous) = 1 - P(none are nervous) - P(both are nervous)

OpenStudy (agent47):

= 1 - 0.82 - 0.0324

OpenStudy (yanasidlinskiy):

0.1476 isn't right.

OpenStudy (yanasidlinskiy):

I don't understand how 0.3276 is the answer.

OpenStudy (agent47):

wait it said at least one,one gives probability of exactly one

OpenStudy (yanasidlinskiy):

I don't understand what you're trying to say.

OpenStudy (agent47):

so P at least one is P exactly one + P exactly two

OpenStudy (sparrow2):